Shift Summary for day shift, 9/17/99

AG (ag@fnal.gov)
Fri, 17 Sep 1999 16:40:05 -0400 (EDT)

Shift Summary for day shift, 9/17/99

On Shift: JCH, AG

Hours of beam available at all intensities: ~ 1h 30min
Hours of beam available at nominal intensity: ~ 0h 45min
Hours of E799 data (special runs not included): 0h 0min
Hours of special runs (laser scans, muon runs...): ~ 1h 0min

Accounting of lost time:

o Accelerator is down
o Difficulties while initializing MUON runs.

========================================================================

Runs ended this shift:

Run 14579 - MUON
Run 14580 - MUON for DC plateaus
Run 145 - BEAM_799, in progress.

Tapes brought to FCC this shift:
(including tapes for online split)

None.

========================================================================

Accesses during shift:

JCH and AG to swap Et board.

========================================================================

Overview of shift:

Beam (of lower intensity and bad quality) comes back at 13:30, when we
take MUON runs and ... beam goes away around 15:00. DC plateaus and TRD HV
scan are partially done, we are waiting for beam.

========================================================================

Description of Shift:

run____spill______time_______comment

------------------------------------------------------------------------

___________________700_____Day shift begins with current estimate for
____________________________beam at 920 hours.

14568____________720_____Start COSMICS run.

___________________800_____Jason L. starts working on bananas, causing
____________________________all sorts of ktevana errors periodically.

___________________825_____New estimate: No estimate.

14569____________831_____Start another COSMICS after Jason L. is done.

___________________850_____Guys are accessing Tevatron.

___________________932_____BobT calls and says that there is going to be
____________________________an access to NM2 as well.

__________________1100_____JCH and AG make an access to swap TE-11 Et
____________________________board. Details in KTeVlog XXI, p.205.

__________________1108_____Tev access complete; Turn-on in progress;
____________________________Estimate 45 min.

14570___________1115_____Page BobT, SashaL, Valeri, Arun, Ashkhan,
____________________________JasonL only. But somehow the rumor spreads
____________________________and control room fills with people very
____________________________quickly.

__________________1120_____Called MCR to ask setup for muon mode.

__________________1134_____Valeri comes in and takes pedestal run to check
____________________________FERA pedestals.

__________________1136_____Turning on the Tevatron.

__________________1200_____Try starting muon run, get 'OCP_RC ERROR'
____________________________during run start.

__________________1220_____MCR calls and promises beam soon. We notice
____________________________that NM2S3 shows ~500amps, although set to 0.
____________________________They'll look at it.

__________________1230_____MCR calls again: NM2SR tripped due to bad card.
____________________________They are trying to replace it.

__________________1253_____The problem with run initialization was that
____________________________disk /ktev3f was not cross mounted and
____________________________BEAM_MUON tries to write to that disk
____________________________remotely. Sydney cross mounted this disk.

14576___________1306____Finally started the muon run. But they are
____________________________still working on NM2S3, so no beam yet.

14576___________1327____Arun stops the run to test config files with
____________________________writing to tapes.

__________________1329_____MCR calls: they fixed NM2SR, we are getting
____________________________muons soon.

__________________1330____And here they come. We are getting muons
___________________________at 3.4e12. We starting muon run...

14579_____1_____1351____Start muon run and get first e799 muons!

14579___________1358____MCR calls: they will be down for 10 minutes.
___________________________We stop the run, update DCGM constants and

14580_____1_____1408____Start muon run for DC plateaus. No beam yet.

14580___________1414____Beam is back. Start doing DC plateaus and TRD HV
___________________________scan.

14580____37_____1458____Beam is gone. P1 magnet problems.

__________________1525____Experts are looking at the problem. DC plateaus
___________________________and TRD HV scan is unfinished.

========================================================================

Follow up Notes on Previous Shift

========================================================================

Items to Follow up on Next Shift

o Ignore VV' low voltage ( V, channel 15 ) SDAQ alarm.